OpenStudy (photon336):

Find the equation of a curve going through point (3,7) and having slope 4x^2-3 at (x,y)

OpenStudy (photon336):

So far this is what I have. \[\frac{ dy }{ dx } = 4x^{2}-3\] \[dy = (4x^{2}-3)~dx\] \[\int\limits dy = \int\limits (4x^{2}-3)~dx\] \[y = \frac{ 4 }{ 3 }x^{3}-3x+C\]

OpenStudy (mathmate):

y=(4/3)x^3-3x+c Substitute (x,y)=(3,7) and solve for C to find the constant.

OpenStudy (math&ing001):

(3,7) to find C

OpenStudy (photon336):

oh wow haha .. that's it I over thought this

OpenStudy (photon336):

Okay, I see now \[\frac{ 4 }{ 3 }(3)^{3}-3(3)+C = 7\]

OpenStudy (photon336):

27+C = 20 So then this would make C = -7

OpenStudy (phi):

don't you mean 27+C = 7 ?

OpenStudy (photon336):

got disconnected

OpenStudy (photon336):

Thanks for clarifying 27+C = 7 So this would make C = -20 => C = -20

OpenStudy (phi):

yes

OpenStudy (mathmate):

Good job @Photon336 !

OpenStudy (photon336):

thank you @mathmate

OpenStudy (mathmate):

You're welcome!